# FALCONEye: Finding Answers and Localizing Content in ONE-hour-long videos with multi-modal LLMs

### v1

@@ 10.2 VLMs Calibration
To measure VLM calibration, we adopt the Reliability Diagrams [12]. These diagrams group all the predictions in bins according to their confidence and measure the gap between confidence and accuracy per each bin. Specifically, we split the $N$ predictions in $M=10$ bins according to their confidence and, for each bin $B_{m}$, we compute its count $N_{m}$ average confidence $C_{m}$ and its average accuracy $A_{m}$. From these diagrams, we may compute the Average Calibration Error (ACE) as,
@@
where $M^{+}$ is the number of non-empty bins [30]. We compute Calibration Count (CC), which quantifies the percentage of predictions above a defined confidence threshold, weighted by their accuracy calibration error (1-CE). For example, the Calibration Count at threshold 0.9 is computed as,
@@
We evaluated various confidence aggregation methods, specifically likelihood, average, and geometric average, as done in similar calibration studies with LLMs [26]. Looking first at the reliability diagrams (Figure 10 (a-f)), we discard likelihood as a suitable confidence metric. The distribution of confidence values is highly spread out, with many answers assigned extremely low confidence. This occurs because likelihood multiplies the probability of all tokens, thus longer answers tend to receive lower confidence scores, making it unreliable for calibration.
@@
Table 11: Calibration metrics comparison with GT 1min-length clips in the open-ended (OQs) questions test split of FALCONBench. Lower values for BS, MCE and ACE, are better.
@@
Figure 10: (a-f) Calibration plots for different probabilities aggregation metrics in the open-question evaluation scenario with the GT 1min clips of the FALCONBench test split. We focus on both LlaVa-Video and Qwen2.5-VL models. (g-l) calibration plots when testing the VLMs with the GT 1min-length clips of the FALCONBench test split for both MCQs (left) and OQs (right) with the geometric average aggregation metric.
@@ 10.3 FALCONEye vs GPT4o
Figure 11 show three example questions and compare the responses from GPT4o and FALCONEye.

### v2

@@ 10.3 VLMs Calibration
To measure VLM calibration, we adopt the Reliability Diagrams [degroot1983comparison]. These diagrams group all the predictions in bins according to their confidence and measure the gap between confidence and accuracy per each bin. Specifically, we split the $N$ predictions in $M=10$ bins according to their confidence and, for each bin $B_{m}$, we compute its count $N_{m}$ average confidence $C_{m}$ and its average accuracy $A_{m}$. From these diagrams, we may compute the Average Calibration Error (ACE) as,
@@
where $M^{+}$ is the number of non-empty bins [neumann2018relaxed]. We compute Calibration Count (CC), which quantifies the percentage of predictions above a defined confidence threshold, weighted by their accuracy calibration error (1-CE). For example, the Calibration Count at threshold 0.9 is computed as,
@@
We evaluated various confidence aggregation methods, specifically likelihood, average, and geometric average, as done in similar calibration studies with LLMs [liu2023litcab]. Looking first at the reliability diagrams (Figure 10 (a-f)), we discard likelihood as a suitable confidence metric. The distribution of confidence values is highly spread out, with many answers assigned extremely low confidence. This occurs because likelihood multiplies the probability of all tokens, thus longer answers tend to receive lower confidence scores, making it unreliable for calibration.
@@
Table 13: Calibration metrics comparison with GT 1min-length clips in the open-ended (OQs) questions test split of FALCON-Bench. Lower values for BS, MCE and ACE, are better.
@@
Figure 10: (a-f) Calibration plots for different probabilities aggregation metrics with the GT 1min clips of the FALCON-Bench test split, and (g-l) calibration plots when testing the VLMs with the GT 1min-length clips of the FALCON-Bench test split.
@@ 10.4 FALCONEye vs GPT4o
Figure 11 shows three FALCON-Bench example questions and compare the responses from GPT4o and FALCONEye.
10.5 FALCON-Bench vs QAEgo4D
The QAEgo4D [Baermann_2022_qaego4d] dataset is the closest to FALCON-Bench in terms of the task definition as it also addresses the problem of VAS, that the authors call question answering visual language grounding (VLG). They also have open ended questions. However, besides the time differences in the video durations –FALCON-Bench with an average duration of $\sim 80$ minutes while QAEgo4D has an average duration of $\sim 8$ minutes–, the type of questions are completely different. QAEgo4D questions are automatically generated from the sparse video narration, focusing on the main object or action: Q: What did I Put in the Pan?A: cheese, or Q: What paint can did I open?A:black paint. Meanwhile, FALCON-Bench questions are human curated to be challenging: Q: What is the number of the train that crosses paths with Lightning McQueen at night?A:A113; Q:What message about the flu appears on a city building?A:Bovril nourishes you to resist ’flu or Q:Which player of Chelsea received a red card?A: Thibaut Courtois.